OpenStudy (anonymous):

(square root of2)/ (1-square root of 5) simplify please help me

OpenStudy (anonymous):

\[\frac{\sqrt{2}}{1-\sqrt{5}}\] \[=\frac{\sqrt{2}}{1-\sqrt{5}}\times \frac{1+\sqrt{5}}{1+\sqrt{5}}\] \[=\frac{\sqrt{2}+\sqrt{10}}{1-5}\]

OpenStudy (anonymous):

aka \[-\frac{\sqrt{2}+\sqrt{5}}{4}\]

OpenStudy (anonymous):

\[\sqrt{2}\div (1-\sqrt{5})\]

OpenStudy (anonymous):

satelilite how do u do fraction??? in equation

OpenStudy (anonymous):

frac{a}{b}